OpenStudy (anonymous):

The position function of a particle is s=t^3-12t, t>and equal to 0. Find the acceleration at the instant when the velocity is 0.

OpenStudy (tylerd):

take the derivative twice

OpenStudy (tylerd):

3t^2-12 = f' or velocity 6t= acceleration

OpenStudy (tylerd):

3t^2=12

OpenStudy (tylerd):

solve for t to find when the velocity is 0

OpenStudy (tylerd):

t^2=4 t=2

OpenStudy (tylerd):

acceleration = 6t plug in 2 for t, acceleration = 12

OpenStudy (anonymous):

oh i see! thanks for your help! I didn't know that we need to figure the t.